OpenStudy (anonymous):

solve for x: x^2+6x-8=0

OpenStudy (anonymous):

It's not factorable, so'll you'll have to complete the square or use quadratic formula.

OpenStudy (anonymous):

The easiest way is to use the quadratic formula \[\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (raden):

yea, It's not factorable but if that question is x^2+6x+8=0 (maybe), can be factored and then take zeros of its factor for get x

OpenStudy (anonymous):

@OD2573 see what to do here?

ganeshie8 (ganeshie8):

maybe q has typo as @RadEn says hmm

OpenStudy (raden):

yea, maybe... hello bos, @mukushla :)